Louisiana State University Health Sciences Center Administration & Finance
 

Friday, June 22, 2007

Dental School :

  • The Central Sterilization demolition has been completed.  The new ducts are being installed, and the walls are laid out.
  • Ceiling grid and tile installation is nearly complete on the 2nd, 3rd, and 4th floors and continues on the 5th, 6th and 7th floors of the Clinic Building. 
    • The installation of the new HVAC flex duct is nearly complete on the 2nd, 3rd and 4th floors and continues on the 5th floor. 
    • Frischhertz has nearly completed reinstalling the lights on the 2nd, 3rd, and 4th floors and continues working on the 5th and 6th floors. 
    • Additional damaged junction boxes and lights were discovered on the 6th and 7th floors and a change order is being written to address the needed repairs. 
    • Firetrol continues to reposition and add sprinkler heads as necessary on the 2nd, 3rd, and 4th floors and continues on the 5th floor.
  • Landis has received the change order to install the new inline zone dampers.  Parts and materials are on order and the installation will begin by the end of next week.
  • The remediation work is nearing completion.
    • We are awaiting area sample results for the 2nd and 3rd floors of the Administration Building.
    • The Air Handling Units 2A and 3A in the Administration Building passed their tests and will be turned on once the area samples come back clean for the corresponding areas. 
    • The Administration Building 4th floor is being cleaned and area samples are scheduled to be taken by close of business today.
    • The cleaning work on the 8th floor of the Clinic is near completion and the area and HVAC viable samples will be taken today or tomorrow.
  • Work on the enhancement of the network infrastructure is progressing.
    • Work is currently under way on the construction of the new network/telecom closets on floors 2 through 8 in the Clinic Building .
      • The build out of each closet includes the installation of plywood back planes, gypsum board and rack installation.
    • The demolition of the old IT closets is underway in the Clinic building. 
    • Cores have been drilled on floors 2 through 8 and sleeves have been installed on floors 2 through 6.

1542 Tulane Ave. :

  • Stratos Elevator started making repairs on Elevator #2 elevator by removing the damaged cab.
  • Ordes Electrical continues installing the 480V 260V AMP service.

Landscaping Project:

  • The installation of edging and planting continue around the Lions/LSU Clinic Building and the CSRB.